Kathy is having an infusion. She needs to pick up her children by 4:30p.m. at the day- care center. If the infusion begins a 10a/m. at 125mL/hour and she needs 750mL , will she be able yo pick up her children on time? Show your work.
Answer #1

If one hour is needed for 125mL, then for 750mL= 750/125 = 6 hours

So, if infusion started at 4:30PM, it will finish by 4:30+6 = 10:30 PM.

Therefore she wouldn't be able to pick up her children on time.